How many grams of h2so4 are needed to react with 29.0 g of nico3
1
Expert's answer
2015-12-02T12:21:08-0500
Solution
Balanced equation is as follows:
H2SO4 + NiCO3 = CO2 + H2O + NiSO4
The amount of NiCO3 given is
n (NiCO3) = m(NiCO3)/M(NiCO3) = 29.0/118.7 = 0.244 mol
The amount of H2SO4 reqired is the same (because of 1:1 reaction stoichiometry), so
n (H2SO4) = 0.244 mol
The mass of H2SO4 reqired is
m(H2SO4) = n(H2SO4)M(H2SO4) = 0.244×98.08 = 23.96≈24 g
Answer: 24 g